OpenStudy (anonymous):

What is the area of a circle with diameter 10 inches?

mathslover (mathslover):

there are two main concepts: 1) area of a circle = \(\large{\pi r^2}\) where r is the radius of the circle 2) diameter of a circle = 2 * radius

mathslover (mathslover):

now .. diameter = 10 inches = 2r r = ?

mathslover (mathslover):

can u tell me what will be r(radius) ?

mathslover (mathslover):

10 = 2* r then r = ? @minecraftlover

OpenStudy (anonymous):

ok i understand :)

mathslover (mathslover):

sure?